QUESTIONS 5A 5B 5C ONLYDue to erratic sales of its sole product-a high-capacity battery for laptop computers- PEM, Inc., has been experiencing finan

ONLY QUESTION 5 PLEASE
Answer #1
Req 5A :
Current unit variable cost = Variable expenses / Units sold = 239400 / 13300 18
Revised unit variable cost = 18 - 3 15
Revised fixed expenses = 177600 + 50000 227600
Contribution margin per unit = Selling price - Unit variable cost = 30 - 15 15
CM ratio = Contribution margin per unit / Selling price = 15 / 30 50%
Break-even point in unit sales = Fixed costs / Contribution margin per unit = 177600 / 15 11840
Break even point in dollar sales = Fixed costs / CM ratio = 177600 / 50% 355200
Req 5B :
Not automated Automated
Total Per unit % Total Per unit %
Sales 612000 30 100% 612000 30 100%
Variable expenses 367200 18 60% 306000 15 50%
Contribution margin 244800 12 40% 306000 15 50%
Fixed expenses 177600 227600
Net operating income 67200 78400
Req 5C :
Answer : Yes
Explanation : If the company sells 20400 units, then the company will earn more net operating income if it automates its operations.